OpenStudy (anonymous):

Convert the numeral to a numeral in base 10 1101 = two

OpenStudy (anonymous):

would it be 1011^20

jimthompson5910 (jim_thompson5910):

1101 base2 = 1*2^3 + 1*2^2 + 0*2^1 + 1*2^0 base 10

OpenStudy (anonymous):

Normally we have numbers in base 10. This means that our digits are 10's, 100's, 1000's, etc. a number like 567 in base 10 can be written as:\[5(10)^2+6(10)^1+7(10)^0\]Now we have base 2. 1011 in base 2 is:\[1(2)^3+0(2)^2+1(2)^1+1(2)^0=?\]so what is it in base 10?

OpenStudy (anonymous):

oh sry, the number in base 2 is 1101. not 1011. the same idea applies though.

OpenStudy (anonymous):

1(10)^3+1(10)^2 +0(10)^1+1(10)^0

OpenStudy (anonymous):

Ah im sry, i should have been a little clearer. To get the number in base 10 when its in another base, you have to calculate:\[1(2)^3+1(2)^2+0(2)^1+1(2)^0=8+4+0+1=13\]So 1101 base 2 is equal to 13 base 10.

OpenStudy (anonymous):

so it only wants me to convert to a numeral in base 10 so my answer would be 13^10

OpenStudy (anonymous):

the ten I wrote should be under the 13 right